OpenStudy (anonymous):

the no. of real solutions of \sin e ^{x}.\cos e^x=2^{x-2}+2^{x+2} is/are?

OpenStudy (nikvist):

do you think \[\sin e^x\cdot\cos e^x=2^{x-2}+2^{x+2}\]

OpenStudy (anonymous):

yea